# Automatically Generating UI Code from Screenshot: A Divide-and-Conquer-Based Approach

### v1

@@ 7. Results & Analysis
7.1. RQ1: Effectiveness of DCGen
This RQ assesses DCGen’s performance in design-to-code generation by comparing it against various direct prompting strategies across multiple metrics.
High-level performance.
Table 4 presents the overall performance of various methods evaluated on the CLIP and BLEU metrics tested on GPT-4o. The results indicate that DCGen achieves the most substantial performance gains in both metrics. Notably, while DCGen and the self-refine approach contribute positively to model performance, the Chain-of-Thought (CoT) prompting approach has a negative impact on the model’s capabilities. Further analysis of the code generated under CoT instruction shows that 70% (84 out of 120) of HTML files contain no CSS content, leading to the generation of purely textual websites devoid of color or style, which indicates the step-by-step visual reasoning procedure impedes the instruction following ability of the model.
Fine-grained metrics.
As detailed in Figure 10, DCGen surpasses both the direct-prompt approach and CoT instruction in fine-grained metrics. DCGen and the self-refine approach enhance GPT-4o’s performance across all metrics, notably in text, color, and block-match metrics, which are indicators of element inclusion and fidelity. Specifically, our method excels in color and block similarity than the self-refine approach. Conversely, while CoT instruction enhances text similarity, it detrimentally affects the model’s ability to reproduce colors, resulting in decreased element fidelity. This observation underscores a significant limitation: current MLLMs struggle with complex visual tasks when left to operate autonomously. The enhanced performance of our method suggests that providing MLLMs with more precise and focused information, such as focused images and fine-grained descriptions, significantly boosts their effectiveness. Figure 11 illustrates two webpage cases (i.e., top and bottom) generated by multiple approaches, demonstrating the superiority of DCGen.
@@ 7.1. RQ1: Effectiveness of DCGen
Figure 10. Fine-grained metrics
Figure 11. Two examples of visual comparison of original and generated websites.
Figure 12. Prompt details for code-base intermediate description for DCGen.
7.2. RQ2: Influence of Language Modality
Drawing inspiration from Chae et al. (Chae et al., 2024; Chen et al., 2022c), which suggests that programming language (PL) may enhance LLMs’ reasoning ability than natural language (NL), we thus investigate which language modality maximizes DCGen’s performance in RQ2.
To answer this RQ, we compare different implementations of DCGen in terms of language modality. In particular, instead of letting MLLMs generate intermediate descriptions for image subdivisions, we directly let the models translate the screenshot segments to code and then perform the assembly process. Prompt details are illustrated in Figure 12. Since UI code has been the intermediate description in DCGen, the assembly MLLM would serve the same role as the generation MLLM and utilize the same prompt.
We investigate the effectiveness of using programming language (PL) versus natural language (NL) in implementing DCGen on two models, GPT-4o and Gemini-1.5-Flash, with the results presented in Table 5. Our results indicate that for GPT-4o, employing PL consistently enhances performance across all visual metrics compared to NL. However, both methods yield similar outcomes for Gemini. The differential performance can potentially be attributed to two factors. Firstly, existing research (Chae et al., 2024; Chen et al., 2022c) suggests that integrating coding into the reasoning process enhances the ability of LLMs to tackle complex tasks. Specifically, by generating accurate code snippets for image segments and subsequently merging them to reconstruct complete website code, MLLMs are more likely to produce precise outputs. Conversely, our observations during the experiments revealed a potential drawback: smaller coding errors tend to propagate and accumulate during the assembly process, more so than with NL descriptions. This error propagation can preserve initial mistakes throughout the assembly process, ultimately compromising the overall quality of the reconstructed website. This phenomenon potentially explains why PL implementation does not always surpass NL implementation.
Answer to RQ2: Although PL implementation can enhance GPT-4o’s performance on all visual metrics, minor coding errors would accumulate in assembly more than with NL descriptions. As a result, the best language modality depends on the specific model.
Table 5. Performance metrics across language modality (%).
Model CLIP BLEU Block Match Position Text Color
GPT4-NL 83.36 8.17 86.54 96.49 84.11 82.13
GPT4-PL 84.93 7.87 88.14 97.53 85.41 83.93
Gemini-NL 77.63 8.81 84.15 96.85 84.30 75.97
Gemini-PL 78.76 8.30 83.81 97.02 81.88 75.58
7.3. Influence of Website Complexity
Recognizing the diverse complexity of real-world websites, this RQ evaluates DCGen’s efficacy on websites with varying complexity levels. We quantify website complexity using the number of HTML tags and focus on changes in three key metrics as functions of website complexity on GPT-4o: CLIP score, BLEU score, and the average of all four fine-grained metrics (Block Match, Text, Position, Color). Figure 15, Figure 15, and Figure 15 illustrate how these metrics vary with website complexity. We observe that DCGen consistently achieves leading performance across all complexity levels for each metric evaluated. Notably, DCGen demonstrates a strong ability to maintain high visual scores (i.e., CLIP and average fine-grain score), indicating its effectiveness in producing code implementations that visually resemble the original webpages, regardless of their complexity. This performance suggests that DCGen is robust to variations in website complexity, effectively handling both simple and intricate website structures.
@@ 7.3. Influence of Website Complexity
Answer to RQ3: DCGen delivers superior performance across various website complexities, indicating its robustness in handling both simple and intricate webpages.
Figure 13. CLIP score versus website complexity.
Figure 14. BLEU score versus website complexity.
Figure 15. Average fine-grain score versus website complexity.
7.4. Generlization of DCGen
@@
In this RQ, we examine the generalizability of DCGen by employing the methodology on different MLLMs as backbones. Tables 7 and 7 display the overall performance of DCGen on these MLLMs. Our results indicate that DCGen is highly adaptable to the Gemini model, achieving notable gains in both visual level and code level metrics compared to other methods. For the Claude-3 platform, DCGen significantly enhances the visual similarity between the original and the generated websites, outperforming competing methods.
@@
Answer to RQ4: DCGen is effective across different MLLMs in generating UI code from design, which demonstrates the generalizability of the proposed divide-and-conquer strategy.
Table 6. Gemini result.
@@
Table 7. Claude-3 result.
@@ 8. Threat to Validity
Limited context length
Current MLLMs have constraints on the length of the context window (e.g., 128K for GPT-4o). Consequently, our approach may fail to handle websites that require a vast number of tokens, exceeding the capacities of MLLMs. Nevertheless, our experiment demonstrates that all of the prompts is below the maximum context length, showing our approach is generalizable for most scenarios.
The selection of backbone models
This study employs three popular instruction-following and multimodal LLMs to demonstrate the effectiveness of DCGen. While other LLMs might be adapted for our methodology, we found that smaller parameter models fail to understand complex prompts. Our future work involves extending our implementation to other emerging models, showcasing DCGen’s generalization abilities.
Unable to handle dynamic websites
Some websites incorporate dynamic features to enhance user interaction, modifying webpage content using server-side scripting languages (e.g., PHP, Python). Although DCGen can only handle static webpages, it is the first segment-aware approach to generate UI code from screenshots, which we believe is an important step toward creating complex dynamic websites.
@@ 9.1. Image-to-code Generation
The most informative AI techniques used for generating code from images can be classified into three broad categories: (i) Convolutional Neural Networks (CNNs) (Asiroglu et al., 2019; Beltramelli, 2018; Xu et al., 2021; Moran et al., 2018a; Chen et al., 2022d; Cizotto et al., 2023; Chen et al., 2018), (ii) Computer Vision (CV) and Optical Character Recognition (OCR) (Nguyen and Csallner, 2015; Natarajan and Csallner, 2018), and (iii) other deep learning models (Wu et al., 2021; Abdelhamid et al., 2020). The work (Beltramelli, 2018) utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). However, DSLs are not widely used in practical development, which inconveniences the industry’s UI developers (Xu et al., 2021). An alternative approach (Chen et al., 2018) using CNNs to extract visual features from UI images is more widely applicable than generating fixed DSL code. Based on CNNs, Andre et al. (Cizotto et al., 2023) pioneered using Class Activation Mapping (CAM), further improving the performance and interpretability of the image-to-code generation task.
@@ 10. Conclusion
In this paper, we first present a motivating study that identifies prevalent failures in MLLMs during the design-to-code generation process. Drawing on these insights, we design and implement DCGen, a novel divide-and-conquer-based framework for effectively generating UI code from web screenshots. We evaluate DCGen across three SOTA MLLMs and demonstrate that it consistently outperforms various direct prompting strategies. Furthermore, DCGen shows robustness against variations in website complexity and is generalizable across different MLLMs, demonstrating its superiority in generating UI code.

### v2

@@ 7. Results & Analysis
7.1. RQ1: How does DCGen compare to various prompting strategies?
This RQ assesses DCGen’s performance in design-to-code generation by comparing it against various prompting strategies across multiple automatic metrics.
7.1.1. High-level Performance
Figure 9 presents the overall performance of various methods evaluated on the CLIP and BLEU metrics tested on GPT-4o. The results indicate that DCGen achieves the most substantial performance gains, outperforming the direct prompting for 13.42% and 1.39% for CLIP and BLEU scores, respectively. Such promising performance demonstrates the effectiveness of our divide-and-conquer approach. Besides, while DCGen and the self-refine approach contribute positively to model performance, the Chain-of-Thought (CoT) prompting approach has a negative impact on the model’s capabilities. A detailed analysis of the code generated under CoT instructions reveals that 70% (84 out of 120) of the HTML files contain no CSS, leading to purely textual websites without color or style. Further investigation suggests that the model, when focused on step-by-step visual reasoning instructions, neglected the directive to include both HTML and CSS in a single output, indicating a limitation in handling large amounts of information simultaneously.
7.1.2. Fine-grained Metrics
[…]
@@ 7.1.2. Fine-grained Metrics
Figure 8. DCGen’s high-level performance on GPT-4o (%).
(a) Fine-grained metrics.
(b) Error rate.
Figure 9. Detailed analysis.
7.2. RQ2: How robust is DCGen for complex designs?
Recognizing the diverse complexity of real-world websites, this RQ evaluates DCGen’s efficacy on websites with varying complexity levels. We quantify website complexity using the number of HTML tags and focus on changes in visual metrics (i.e., CLIP score and the average of all four fine-grained metrics) using GPT-4o as the backbone. In our sample dataset, the number of HTML tags varies from 24 to 298, where a higher number of tags indicates increased complexity in the websites.
Figure 10(a), and Figure 10(b) showcase DCGen’s performance across varying degrees of complexity. DCGen consistently achieves leading performance across all metrics, demonstrating its strong ability to maintain high visual scores over varying complexity levels. Specifically, as website complexity escalates from 24 tags to 298 tags, DCGen’s performance remains stable, with only a minor 5.1% decline in effectiveness as measured by the CLIP score. Even though the division stage is straightforward, such consistent performance underlines the robustness and effectiveness of DCGen in handling website complexity.
@@ 7.2. RQ2: How robust is DCGen for complex designs?
Answer to RQ2: DCGen delivers superior performance across various website complexities, indicating its robustness in handling both simple and intricate webpages.
(a) CLIP score vs. website complexity.
(b) Avg. fine-grain score vs. website complexity.
Figure 10. DCGen’s robustness for complex websites.
7.3. RQ3: How generalizable is DCGen for other SOTA MLLMs?
@@
In this RQ, we examine the generalizability of DCGen by employing the methodology on different MLLMs as backbones. Tables 5 and 5 display the overall performance of DCGen on Gemini-1.5 and Claude-3. Our results indicate that DCGen is highly adaptable to the Gemini model, achieving notable gains of 3.9% in visual level metrics compared to other methods since in general, the decomposition allows models to conduct more focus on smaller visual segments and more manageable sub-generation tasks. For the Claude-3, DCGen also significantly enhances the visual similarity between the original and the generated websites by 3.8%, outperforming competing methods. As a result, the experiment demonstrates the generalizability of DCGen on various MLLMs, showing the superiority of the divide-and-conquer strategy on UI Code generation.
@@
Answer to RQ3: DCGen is effective across different MLLMs in generating UI code from design, which demonstrates the generalizability of the proposed divide-and-conquer strategy.
Table 4. DCGen’s performance on Gemini-1.5.
@@
Table 5. DCGen’s performance on Claude-3.
@@
7.4. RQ4: What is the human’s feedback on DCGen in real-world application development workflows?
While automatic metrics offer a detailed breakdown of model performance, it is crucial to consider human feedback to validate the effectiveness of DCGen. To this end, we conduct a series of human evaluations to compare the performance of different methods and directly assess the practical utility of DCGen in real-world application development workflows.
7.4.1. Pairwise Method Comparison
In this study, we recruit five PhD students from our university as annotators. Following the conventional practice of evaluating instruction following LLMs (Zhou et al., 2023; Dubois et al., 2023), we ask the annotators to compare pairs of generated webpages (one from the baseline, the other from the tested methods) to decide which one is more similar to the reference. We use GPT-4o Direct Prompting as the baseline and collect the other three methods’ Win/Tie/Lose rates against this baseline. Each pair will count as Win (Lose) only when Win (Lose) receives the majority vote ($\geq 3$). All other cases are considered Tie.
Results. We conducted pairwise comparisons between DCGen and baseline methods, with the results shown in Figure 11. DCGen outperformed direct prompting in 62% of cases and underperformed in just 8%, showing its effectiveness in generating UI code. Meanwhile, we find that DCGen is substantially better than other baselines, while both self-revision and CoT prompting show marginal improvements over direct prompting (winning in 38-40% cases and losing in 16-18% cases). These findings suggest that automatic metrics align with human perception, demonstrating that DCGen can produce more visually similar target webpages than competing methods under human evaluation.
Figure 11. Human pairwise preference evaluation results with GPT-4o Direct Prompting as the baseline. A higher win rate and lower lose rate suggest better quality as judged by human annotators.
7.4.2. Usefulness Evaluation
While the pairwise evaluation offers a visual comparison among methods, we further assess the practicality of DCGen in accelerating the development workflow. In specific, we implement DCGen into a user-friendly demo tool (detailed in Section 8) and recruit four PhD students and research staff from our institution to participate in our study, following the methodology from Chen et al. (Chen et al., 2018). Each participant is tasked with converting two website UI images into corresponding UI code. The two images are selected to represent different levels of complexity: one simple and one more complex. Participants are randomly assigned to one of two groups: the experimental group (P1, P2), which uses DCGen to implement the UI code, and the control group (P3, P4), which is allowed to use any preferred tool (e.g., GPT-4o). A pre-study background survey confirms that all participants have similar development experience, being familiar with HTML and CSS and having previously developed one to two web applications for their work. Each participant has up to 20 minutes to complete each design.
We record the time each participant takes to implement the UI designs and capture screen recordings of their development processes, which are available in our supplementary materials. After the experiment, two research staff, who are not involved in the study, evaluate the similarity of the implemented GUIs to the original UI images using a five-point Likert scale (Joshi et al., 2015) (1: not similar at all, 5: almost identical). The judges are blinded to which group produces each GUI.
Results. Figure 12 highlights UI implementation results from the experiment group (w/ DCGen) and control groups (w/o DCGen). First, the box plot in Figure 12(a) shows that the experiment group completes the design-to-code task twice as fast as the group without DCGen. The control group also shows greater variability, with some participants taking up to 20 minutes. In fact, the average time of the control group is underestimated, because one participant fail to complete at least one UI image within 20 minutes, which means that the participant needs more time in the real development. In contrast, all participants in the experiment group finish all the tasks within 15 minutes. Second, as shown in figure 12(b), the experiment group consistently achieves higher similarity scores compared to the control group. The Pearson Correlation Coefficient coefficient between the similarity scores given by the two judges in 12(b) is 0.71, suggesting a strong agreement between the judges. Finally, figure 12(c) summarizes these results, confirming that the experiment group using DCGen has lower completion times and higher similarity scores on average, indicating our tool can significantly help the experimental group implement webpages faster and more similar to the UI design images.
(a) Time distribution (minutes).
(b) Similarity distribution.
Measures Ctrl Exp
Time (min) 14.5 8
Similarity 2.25 3.75
(c) Average result.
Figure 12. The comparison of the experiment and control group on time and similarity.
Answer to RQ4: Human evaluations demonstrate that DCGen can produce more visually similar target webpages than competing methods and the DCGen tool can significantly help developers implement webpages faster and more similar to the UI designs.
8. Discussion
Based on the users’ experience of interacting with DCGen, we discuss how it could contribute to transferring UI designs to code for developers, as well as identify potential limitations of DCGen for future improvements.
Figure 13. Success cases. The first row shows an example from RQ1, and the rest shows examples from the usefulness study (RQ4).
Success stories. The first row of Figure 13 illustrates a success case from Section 7 RQ1, which shows that compared to the self-refine method, DCGen restores the color and layout of the webpage, and removes nonexistent pictures from the webpage thus reducing distortion. The rest of Figure 13 provides examples from the usefulness study (RQ4). DCGen demonstrates its effectiveness by significantly reducing element omissions and distortions, leading to faster development and improved webpage quality.
In the usefulness study (RQ4), we observe that both developers in the control group follow a similar iterative refinement pattern to generate UI code, even though they are unaware of the DCGen’s methodology. Specifically, they start by prompting the model to produce an initial webpage, then iteratively refine the unsatisfactory parts by pointing out current errors and asking MLLMs for improvements in follow-up conversations. This process repeats until they are satisfied with the results or the time runs out. Meanwhile, developers also fix missing elements issues by providing more detailed prompts with a screenshot of the specific areas. Such manual narrowing of task scope aligns with DCGen’s automated divide-and-conquer approach.
Tool features. Figure 14 shows a screenshot of DCGen’s user interface as a development tool, with the original image on the left and the code and rendered webpage on the right. DCGen enables a lot of features, allowing developers to generate an entire webpage (Figure 14(a)), inspect the code for individual image segments (Figure 14(b)), and regenerate code for any specific segment (Figure 14(c)). By automating the divide-and-conquer process, DCGen eliminates the need for detailed, repetitive prompts, speeding up development and enabling the creation of higher-quality webpages in less time compared to the control group.
(a) Generate code for the entire webpage.
(b) View the code of an image segment.
(c) Generate code for an image segment.
Figure 14. DCGen user interface and functionalities.
Failure analysis. While DCGen effectively reduces element omission and distortion, its improvement in addressing element misarrangement is limited. A failure case is illustrated in Figure 14(a), within the orange rectangular area. In the original image, the paragraphs are horizontally arranged, but the generated paragraphs are vertically aligned. Upon further investigation, we found that this error occurs during the assembly process, as the leaf solver LLM correctly generates horizontally arranged paragraphs for this area, as shown in Figure 14(c). We attribute the issue to the inherent visual limitations of MLLMs, as identified by Tong et al. (Tong et al., 2024), which suggest that MLLMs still face challenges in recognizing certain simple visual patterns, such as positional arrangements.
However, since DCGen allows users to review and regenerate code for any specific image segment, the erroneous code can be easily replaced with the correct code for the corresponding image segment in the final webpage. Thus, this failure case has minimal impact on DCGen’s overall practical effectiveness.
@@ 9. Threat to Validity
Limited context length. Current MLLMs have constraints on the length of the context window (e.g., 128K for GPT-4o). Consequently, our approach may fail to handle websites that require a vast number of tokens, exceeding the capacities of MLLMs. Nevertheless, our experiment demonstrates that all of the prompts is below the maximum context length, showing our approach is generalizable for most scenarios.
The selection of backbone models. This study employs three popular instruction-following and multimodal LLMs to demonstrate the effectiveness of DCGen. While other LLMs might be adapted for our methodology, we found that smaller parameter models fail to understand complex prompts. Our future work involves extending our implementation to other emerging models, showcasing DCGen’s generalization abilities.
Robustness of set thresholds and chosen separation strategy. Assumptions regarding the segmentation strategy and manually set thresholds may limit DCGen’s applicability in more complex designs. To address this concern, we conduct RQ2 to evaluate DCGen’s robustness to complex designs. The results show that DCGen delivers superior performance across various website complexities, indicating its robustness in handling both simple and intricate webpages.
@@ 10.1. Image-to-code Generation
The most informative AI techniques used for generating code from images can be classified into three broad categories: (i) Convolutional Neural Networks (CNNs) (Asiroglu et al., 2019; Beltramelli, 2018; Xu et al., 2021; Moran et al., 2018a; Chen et al., 2022c; Cizotto et al., 2023; Chen et al., 2018), (ii) Computer Vision (CV) and Optical Character Recognition (OCR) (Nguyen and Csallner, 2015; Natarajan and Csallner, 2018), and (iii) other deep learning models (Wu et al., 2021; Abdelhamid et al., 2020). The work (Beltramelli, 2018) utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). However, DSLs are not widely used in practical development, which inconveniences the industry’s UI developers (Xu et al., 2021). An alternative approach (Chen et al., 2018) using CNNs to extract visual features from UI images is more widely applicable than generating fixed DSL code. Based on CNNs, Andre et al. (Cizotto et al., 2023) pioneered using Class Activation Mapping (CAM), further improving the performance and interpretability of the image-to-code generation task.
@@ 11. Conclusion
In this paper, we first present a motivating study that identifies prevalent failures in MLLMs during the design-to-code generation process. Drawing on these insights, we design and implement DCGen, a novel divide-and-conquer-based framework for effectively generating UI code from web screenshots. We evaluate DCGen across three SOTA MLLMs and demonstrate that it consistently outperforms various prompting strategies. Furthermore, DCGen shows robustness against variations in website complexity and is generalizable across different MLLMs, demonstrating its superiority in generating UI code.